OpenStudy (anonymous):

Chemistry terms help? I last took chemistry in the ninth grade, and I learned all the terms in Turkish- could someone help me express this ( 2NO (g) + 5H2 (g) --> 2NH3 (g) + 2H2O (g) ) equation in words? And explain to me what means what, etc?

OpenStudy (aaronq):

so this reads as "two molecules nitric oxide in the gas phase react with 5 molecules of hydrogen gas to produce 2 molecules of ammonia gas and 2 molecules of water in the gas phase". What i did was repetitive but i wanted the information written in the equation to be clear. How the information is conveyed is described below: \(\Large \sf \color{red}2~NO_{ (g)} + \color{red}5~H_{\color{blue}{2}~ (g)}\rightarrow \color{red}2~NH_{\color{blue}{3} (g)} + \color{red}2~H_{\color{blue}2}O_{ (g)}\) On the left of the arrow we have reactants and on the right of the arrow, products. The chemical formula indicates the types of molecules reacting. The subscripts (in \(\sf \color{blue}{blue}\)) are parts of the chemical formula and indicate how many of each type of atom there are in a molecule. \(\sf \Large H_{\color{blue}2 }O\) is composed of 2 Hydrogen atoms and 1 Oxygen atom. (when no subscript is written (like in Oxygen above) it's assumed to be 1) Other Symbols indicate the phase, of that chemical species: (g) gas, (l) liquid, (aq) aqueous, (s) solid. In red, \(\sf (stoichiometric)~\color{red}{coefficient}\) indicates how many molecules are reacting - it follows the distributive property (from math), e.g. \(\sf \Large \color{red} 3~H_{\color{blue}2}O\rightarrow \color{red} 3*\color{blue}2~H+\color{red}3*O=6~H+3~O\) \(\sf \Large \color{red} 2~(2x+3y)=\color{red}2*2~x+\color{red}2*3~y=4~x+6~y\) Let me know if you have further questions.